# Arc-length characterization of finite, radial growth patterns

### v1

@@ 5 Using the telegraph signal to find the centre
The choice of the location of the centre of a roughly radially symmetric pattern is to some extent arbitrary. The simplest method to find the centre of mass of a pattern is to locate its center of mass. However, for the colloidal drying patterns, the centre of mass is skewed by the transversal thick fingers forming at the rim of the pattern. These thick fingers tend to form unevenly, hence the centre of mass is clearly different from any visually selected centre. If the pattern is a fully connected network, like a DLA, it is possible to follow the fingers towards their centre. Following the idea of Horton and Strahlers works on complexity in river networks 5, 12, we may skeletonize a pattern and count the numbers of bifurcations at each node in the skeleton network. However, this method may provide results that are clearly off the centre, as it a) requires a fully contiguous pattern and b) cause arbitrariness where there are multiple equally ordered points.
@@
$\mathcal{T}_{S}(\theta;X_{i})=\frac{1}{r_{\text{max}}-r_{\text{min}}}\int_{r_{\text{min}}}^{r_{\text{max}}}\mathcal{T}(r,\theta;X_{i})dr.$
We then calculate the power spectrum of the signal as $G(\hat{\theta})=\int_{0}^{\infty}|\mathcal{T}_{S}(\theta;X_{i})|^{2}d\theta,$ where $\hat{\theta}$ denotes the Fourier mode of the angle. A clear periodicity in the summed telegraph signal will result in a sharp peak in the power spectrum. We therefore use the power spectrum to define an information entropy, as $S(X_{i})=\int_{0}^{\infty}G(\hat{\theta};X_{i})\ln{G(\hat{\theta};X_{i})}d\hat{\theta}$. The best centre, based on our information entropy, is simply picked as the extremal point $X_{c}=\max(S(X_{i}))$. Figure 7a-c shows a comparison of two potential centres for the RBP figure, along with their power spectra. Summing the spectra as an entropy, we obtain the map in Figure 7d. In Figure 7e the comparison between the maximum entropy, centre of mass and Horton-Strahler approaches shows that our maximum entropy determination of the figure’s centre is the most accurate.
@@
Figure 8: (a) Resulting entropy map of patterns P1 and P2, using the method presented in this paper. (b) Resulting centre points, compared with the two alternative methods.
@@ 6 Conclusions
We find that the colloidal drying patterns P1-P3 fall somewhere between the characteristic DLA and Random Birthday Pie patterns. Most importantly, we find the competition between the generation and merging of fingers sustains a fairly stable number of fingers, up to a point where the fingers merge more rapidly. Finally, we use the arc lengths to map an information entropy on the pattern, which we may use to pick a centre where this is not trivial.
The method can be applied to patterns that are not simply radially invariant or self-similar, but fall somewhere between the two.
@@
This work was supported by the Research Council of Norway through its Center of Excellence funding scheme, project number RCN 262644. The authors would like to thank Daan Frenkel for fruitful discussions and extended help in writing this manuscript.

### v2

@@ 6 Using the telegraph signal to find the centre
The choice of the location of the centre of a roughly radially symmetric pattern is to some extent arbitrary. The simplest method to find the centre of mass of a pattern is to locate its center of mass. However, for the colloidal drying patterns, the centre of mass is skewed by the transversal thick fingers forming at the rim of the pattern. These thick fingers tend to form unevenly, hence the centre of mass is clearly different from any visually selected centre. If the pattern is a fully connected network, like a DLA, it is possible to follow the fingers towards their centre. Following the idea of Horton and Strahlers works on complexity in river networks 7, 15, we may skeletonize a pattern and count the numbers of bifurcations at each node in the skeleton network. However, this method may provide results that are clearly off the centre, as it a) requires a fully contiguous pattern and b) cause arbitrariness where there are multiple equally ordered points.
@@
$\mathcal{T}_{S}(\theta;X_{i})=\frac{1}{r_{\text{max}}-r_{\text{min}}}\int_{r_{\text{min}}}^{r_{\text{max}}}\mathcal{T}(r,\theta;X_{i})\mathrm{d}r.$
We then calculate the power spectrum of the signal as $G(\hat{\theta})=\int_{0}^{\infty}|\mathcal{T}_{S}(\theta;X_{i})|^{2}\mathrm{d}\theta,$ where $\hat{\theta}$ denotes the Fourier mode of the angle. A clear periodicity in the summed telegraph signal will result in a sharp peak in the power spectrum. We therefore use the power spectrum to define an information entropy, as $S(X_{i})=\int_{0}^{\infty}G(\hat{\theta};X_{i})\ln{G(\hat{\theta};X_{i})}\mathrm{d}\hat{\theta}$. The best centre, based on our information entropy, is simply picked as the extremal point $X_{c}=\max(S(X_{i}))$. Figure 10a-c shows a comparison of two potential centres for the RBP figure, along with their power spectra. Summing the spectra as an entropy, we obtain the map in Figure 10d. In Figure 10e the comparison between the maximum entropy, centre of mass and Horton-Strahler approaches shows that our maximum entropy determination of the figure’s centre is the most accurate.
@@
Figure 11: (a) Resulting entropy map of patterns P1 and P2, using the method presented in this paper. Green in the scale bar corresponds to the best location of the centre. (b) Resulting centre points, compared with the two alternative methods.
@@ 7 Conclusions
We find that the colloidal drying patterns P1-P3 fall somewhere between the characteristic DLA and Random Birthday Pie patterns. Most importantly, we find the competition between the generation and merging of fingers sustains a fairly stable number of fingers, up to a point where the fingers merge more rapidly. We compute the correlation between the telegraph signals, and relate them to a characteristic exponential decay with a persistence length of $\xi=0.063$ mm. With a few assumptions, we are also able to relate the information from the arc-lengths to describe how the pattern deposits its colloidal mass and to the colloid concentration of the droplet, as it evaporates. This is also used to characterize the fractality of the patterns, showing the fractal dimension is similar to the one of DLA. Finally, we use the arc lengths to map an information entropy on the pattern, which we may use to pick a centre where this is not trivial.
Hence, the method can be applied to patterns that are not simply radially invariant or self-similar, but fall somewhere between the two.
8 Data Availability
All the images, methods and Python codes used in this paper is available in a public Zenodo repository 5. This includes the method used to generate the synthetic patterns (Diffusion-Limited Aggregation and Random Birthday Pies), the method we used to binarize our experimental images and the full, annotated Python codes we developed to perform our analyses.
@@ 8 Data Availability
This work was supported by the Research Council of Norway through its Center of Excellence funding scheme, project number RCN 262644. The authors would like to thank Daan Frenkel and Benjamin Rotenberg for fruitful discussions and critically reading this manuscript.